statistics show that 80% of crimes are never solved.suppose a neighborhood district in a large city suffers repeated crimes the police are investigating 6 property crime cases in this district.
a) what is the probability that none will ever be solved
b0probability at least one will be solved
c0expected # of crimes that will be solved and what is the standard deviation

Answer by stanbon(75887)

statistics show that 80% of crimes are never solved.suppose a neighborhood district in a large city suffers repeated crimes the police are investigating 6 property crime cases in this district.
a) what is the probability that none will ever be solved
0.8^6 = 0.2621
---------------------------
b) probability at least one will be solved
P(at least one solved) = 1 - P(none solved) = 1-0.2621 = 0.737856...
--------------------------------------
c) expected # of crimes that will be solved and what is the standard deviation
--------------
Expected number = np
Expected number = 6*0.8 = 4.8
Std = sqrt(6*0.8*0.2) = 0.9798
======================================
Cheers,
Stan H.